Arun bought a pair of skates at a sale where the discount given was 20%. If the amount he pays is ₹ 1,600, find the marked price.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Discount percent = Discount × 100
Marked price
20 =Discount× 100
x
Discount =20/100 × x
=1/5𝑥
Also,
Discount = Marked price − Sale price
1/5𝑥 = 𝑥 − ₹1600
x –1/5x =₹1600
4/5𝑥 =₹1600
𝑥 =₹1600×5/4
= 2000
The marked price was₹2000

Thanks